# Numerical boundary flux functions that give provable bounds for nonlinear initial boundary value problems with open boundaries

### v1

@@ 1 Introduction
Discontinuous Galerkin Spectral Element Methods (DGSEMs) are implemented in numerous packages [1, 2, 3, 4, 5]. The methods are attractive because they feature geometric flexibility with high order accuracy. They can also be stabilized by bounding some measure of the solution (e.g. entropy [6, 7] and/or kinetic energy [8, 9]) provided that a split form approximation is used in the interior, and that the boundary conditions are implemented stably, e.g. [10, 11, 12]. True energy stability, which requires a bound on all dependent variables, has not been previously shown except when the mathematical entropy is also an energy.
@@
Here, we derive and prove a condition on numerical boundary fluxes for inflow-outflow boundaries that ensures that the mathematical entropy rate is bounded solely by boundary data. We then derive explicit examples of such flux functions that satisfy these conditions for the Burgers equation and for the shallow water equations, for which entropy boundedness is also energy stability. Finally, we present numerical examples using a DGSEM that show that standard flux functions may work, but can fail, whereas the new fluxes do not.
@@ 2 Brief overview of entropy analysis
where $\mathbf{q}$ is the state vector of conserved variables and $\mathbf{f}$ is the flux. The system is solved in a domain $\Omega$ with initial conditions $\mathbf{q}(x,0)=\mathbf{q}_{0}(x)$ and boundary conditions $\mathbf{q}(x,t)=\mathbf{g}(t)$ in $\partial\Omega$.
@@
To weakly incorporate the boundary conditions into the entropy evolution, we introduce a modified flux that is a function of the internal solution state and the external data, $\mathbf{f}^{*}(\mathbf{g},\mathbf{q})$, to have
$\int\limits_{\Omega}S_{t}\,\mathrm{d}\Omega+\int\limits_{\partial\Omega}\mathbf{v}^{T}\mathbf{f}^{*}(\mathbf{g},\mathbf{q})\,\operatorname{dS}-\int\limits_{\Omega}\mathbf{v}_{x}^{T}\mathbf{f}\,\mathrm{d}\Omega=0.$ (2.4)
@@
$\int\limits_{\Omega}S_{t}\,\mathrm{d}\Omega+\int\limits_{\Omega}\mathbf{v}^{T}\mathbf{f}_{x}\,\mathrm{d}\Omega+\int\limits_{\partial\Omega}\mathbf{v}^{T}\left(\mathbf{f}^{*}(\mathbf{g},\mathbf{q})-\mathbf{f}\right)\,\operatorname{dS}=0.$ (2.5)
@@
$\int\limits_{\Omega}S_{t}\,\mathrm{d}\Omega+\int\limits_{\partial\Omega}\left\{f^{\varepsilon}+\mathbf{v}^{T}\left(\mathbf{f}^{*}(\mathbf{g},\mathbf{q})-\mathbf{f}\right)\right\}\,\operatorname{dS}=0.$ (2.6)
@@
$\frac{d\mathcal{E}}{dt}+\int\limits_{\partial\Omega}\left\{f^{\varepsilon}+\mathbf{v}^{T}\left(\mathbf{f}^{*}(\mathbf{g},\mathbf{q})-\mathbf{f}\right)\right\}\,\operatorname{dS}=0.$ (2.7)
@@ 3 The split form discontinuous Galerkin spectral element method
There is a one-to-one discrete analog comparing (3.1) with the continuous result (2.7). The entropy is bounded if the integrand along physical boundaries in (3.1) is bounded by data, and the only freedom is in the choice of the numerical boundary flux. For open boundaries we know of no provably bounded nonlinear boundary numerical flux functions. It is not obvious how to discern directly how the entropy flux, $F_{n}^{\varepsilon}$, can correctly “balance” with the numerical flux penalty, $\mathbf{F}^{*}_{n}-\mathbf{F}_{n}$, to ensure that the boundary integral is bounded by data. To find such a balance is the topic of this paper.
@@ 4 A condition for nonlinearly bounded boundary flux functions
We derive a condition on the flux function $\mathbf{F}^{*}$ in (3.1) to guarantee entropy boundedness by collecting results from a series of papers on nonlinearly stable boundary conditions found in [25, 26, 27, 28]. The idea is to write the entropy flux, $F_{n}^{\varepsilon}$, in terms of a boundary matrix $\underline{\textnormal{A}}$ that can be diagonalized by a transformation $\underline{\textnormal{A}}=\underline{\textnormal{T}}\,\underline{\textnormal{$\Lambda$}}\,\underline{\textnormal{T}}^{T}$ so that
@@
where $\mathbf{W}=\underline{\textnormal{T}}^{T}\mathbf{U}$ defines a vector of (nonlinear) characteristic variables. The sign of the terms in the diagonal matrix $\underline{\textnormal{$\Lambda$}}$ specifies the minimal number of boundary conditions required to obtain a bound on the solution in terms of external data [31, 25, 32, 26]. For linear problems, $\underline{\textnormal{T}}$ is the matrix of eigenvectors and $\underline{\textnormal{$\Lambda$}}$ is the matrix of eigenvalues of the matrix $\underline{\textnormal{A}}$. For nonlinear problems we will use a congruence transformation as discussed below.
The decomposition in (4.1) exposes how solution information propagates in the directions determined by the sign of the terms in the diagonal matrix $\underline{\textnormal{$\Lambda$}}$, so we decompose $\underline{\textnormal{$\Lambda$}}=\underline{\textnormal{$\Lambda$}}^{+}+\underline{\textnormal{$\Lambda$}}^{-}$, where $\underline{\textnormal{$\Lambda$}}^{\pm}=\frac{1}{2}\left(\underline{\textnormal{$\Lambda$}}\pm|\underline{\textnormal{$\Lambda$}}|\right)$ into matrices with only positive or negative entries. We associate outgoing and incoming (nonlinear) characteristic variables with those positive and negative terms by
@@
Again, following [25, 26, 27], we prove the following theorem:
@@
$\mathbf{W}^{T}\underline{\textnormal{$\Lambda$}}\mathbf{W}+\mathbf{U}^{T}\left(2\underline{\textnormal{T}}\,\underline{\textnormal{I}}^{-}\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\left(\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\mathbf{W}^{-}-\mathbf{G}\right)\right)\geq-\mathbf{G}^{T}\mathbf{G}.$ (4.3)
@@
From Thm. 4.1, we have,
@@
We compare the terms in (4.8) with (4.3) and (4.6) to arrive at the desired result. ∎
@@ 5 A stable nonlinear inflow boundary flux for the Burgers equation
To illustrate the process and to motivate and highlight the vital steps, we demonstrate the use of Thm. 4.2 by deriving an energy stable boundary flux for the Burgers equation. The Burgers equation, although nonlinear, is simple enough to derive the boundary flux function in (3.1) from first principles. We show here that the first principle derivation gives the same result as Thm. 4.2.
5.1 A derivation using first principles
@@
When $u>0$, $\left.\frac{1}{3}u^{3}\right|_{x_{R}}>0$ and has the correct sign to dissipate energy, so no boundary condition is needed at $x_{R}$. At the left boundary, if we ensure that the contribution is a square of the specified data, i.e., if the contribution from the left (inflow) side is equal to some $G^{2}$, where $G(t)$ is specified data, then the energy is bounded exclusively by data, i.e., $\mathcal{E}_{t}\leq G^{2}$ and $\mathcal{E}(T)\leq\int_{0}^{T}G^{2}dt$.
@@
$\overline{W}=G=\sqrt{\frac{1}{3}|u_{\textrm{ext}}|}u_{\textrm{ext}},$ (5.4)
then $BT\leq G^{2}$. This means that, unlike in the linear analysis, one does not specify the solution at the boundary, but (5.4) instead, which includes the wave speed.
Since (3.1) imposes the boundary conditions in weak form as a penalty, we need the appropriate penalty that ensures that $\mathcal{E}$ is bounded by data in the same way. From (5.4), we penalize by adding a term proportional to $\overline{W}-G$,
@@
where $l(x)$ is a lifting operator defined as $\int_{x_{L}}^{x_{R}}\phi l(\psi)dx=\phi\psi|_{x_{L}}$ [33, 34]. The penalty is on the characteristic variable $\overline{W}$, which, again, differs from linear problems where the dependent variable $u$ is penalized towards the external solution, $u_{ext}$.
@@
When we substitute for $\overline{W}$, $G$, $F$, and $\sigma$, use the fact that $u>0$, and rearrange, we find the numerical boundary flux function that satisfies Thm. 4.2 and ensures that the entropy is bounded by data to be
@@
This boundary term is not a square bounded by data independent of $u$, and hence do not lead to an estimate. If, instead, we use the local Lax-Friedrichs (LLF) flux [35] at the boundary, then
$BT=\left\{\begin{aligned} &-\frac{u^{3}}{6}+\frac{uu_{ext}^{2}}{2}-\frac{u(u-u_{ext})^{2}}{4},\quad u>u_{ext}\\[5.69046pt] &-\frac{u^{3}}{6}+\frac{uu_{ext}^{2}}{2}+\frac{u(u-u_{ext})^{2}}{4},\quad u<u_{ext}.\end{aligned}\right$. (5.13)
Again, the boundary term is not bounded independently of the interior state, $u$, and using this flux does not lead to an entropy bound.

### v2

# arXiv:2511.04197v2[math.NA] 02 Mar 2026

@@ 1 Introduction
Discontinuous Galerkin Spectral Element Methods (DGSEMs) are implemented in numerous packages [1, 2, 3, 4, 5]. The methods are attractive because they feature geometric flexibility with high order accuracy. They can also be stabilized by bounding some measure of the solution (e.g. entropy [6, 7] and/or kinetic energy [8, 9]) provided that a split form approximation is used in the interior, and that the boundary conditions are implemented stably, e.g. [10, 11, 12]. True energy stability, which requires a bound on all dependent variables, has not been previously shown except when the mathematical entropy is also an energy [13, 14], otherwise additional positivity assumptions as in [6, 7, 15] are necessary.
@@
Here, we derive and prove a condition on numerical boundary fluxes for inflow-outflow boundaries that ensures that the mathematical entropy rate is bounded solely by boundary data without additional assumptions on solution quantities. We then derive explicit examples of such flux functions that satisfy these conditions for the Burgers equation and for the shallow water equations, for which entropy boundedness is also energy stability. Finally, we present numerical examples using a DGSEM that show that standard flux functions may work, but can fail, whereas the new fluxes do not.
@@ 2 Brief overview of entropy analysis
where $\mathbf{q}$ is the state vector of conserved variables and $\mathbf{f}$ is the flux. The system is solved in a domain $\Omega$ with initial conditions $\mathbf{q}(x,0)=\mathbf{q}_{0}(x)$ and boundary conditions $\mathbf{q}=\mathbf{q}_{\text{ext}}(x,t)$ in $\partial\Omega$.
@@
To weakly incorporate the boundary conditions into the entropy evolution, we introduce a modified flux that is a function of the internal solution state and the external data, $\mathbf{f}^{*}(\mathbf{q},\mathbf{q}_{\text{ext}})$, to have
$\int\limits_{\Omega}S_{t}\,\mathrm{d}\Omega+\int\limits_{\partial\Omega}\mathbf{v}^{T}\mathbf{f}^{*}(\mathbf{q},\mathbf{q}_{\text{ext}})\,\operatorname{dS}-\int\limits_{\Omega}\mathbf{v}_{x}^{T}\mathbf{f}\,\mathrm{d}\Omega=0.$ (2.4)
@@
$\int\limits_{\Omega}S_{t}\,\mathrm{d}\Omega+\int\limits_{\Omega}\mathbf{v}^{T}\mathbf{f}_{x}\,\mathrm{d}\Omega+\int\limits_{\partial\Omega}\mathbf{v}^{T}\left(\mathbf{f}^{*}(\mathbf{q},\mathbf{q}_{\text{ext}})-\mathbf{f}\right)\,\operatorname{dS}=0.$ (2.5)
@@
$\int\limits_{\Omega}S_{t}\,\mathrm{d}\Omega+\int\limits_{\partial\Omega}\left\{f^{\varepsilon}+\mathbf{v}^{T}\left(\mathbf{f}^{*}(\mathbf{q},\mathbf{q}_{\text{ext}})-\mathbf{f}\right)\right\}\,\operatorname{dS}=0.$ (2.6)
@@
$\frac{d\mathcal{E}}{dt}+\int\limits_{\partial\Omega}\left\{f^{\varepsilon}+\mathbf{v}^{T}\left(\mathbf{f}^{*}(\mathbf{q},\mathbf{q}_{\text{ext}})-\mathbf{f}\right)\right\}\,\operatorname{dS}=0.$ (2.7)
@@ 3 The split form discontinuous Galerkin spectral element method
There is a one-to-one discrete analog comparing (3.1) with the continuous result (2.7). The entropy is bounded if the integrand along physical boundaries in (3.1) is bounded by data, and the only freedom is in the choice of the numerical boundary flux. As mentioned in the introduction, for open boundaries, provably bounded nonlinear numerical flux functions for the shallow water equations without additional stabilizing assumptions are unavailable. It is not obvious how to discern directly how the entropy flux, $F_{n}^{\varepsilon}$, can correctly “balance” with the numerical flux penalty, $\mathbf{F}^{*}_{n}-\mathbf{F}_{n}$, to ensure that the boundary integral is bounded by data. To find such a balance is the topic of this paper.
@@ 4 A condition for nonlinearly bounded boundary flux functions
We derive a condition on the flux function $\mathbf{F}^{*}$ in (3.1) to guarantee entropy boundedness by collecting results from a series of papers on nonlinearly stable boundary conditions found in [13, 28, 29, 30]. The idea is to write the entropy flux, $F_{n}^{\varepsilon}$, in terms of a symmetric boundary matrix $\underline{\textnormal{A}}$ that can be diagonalized by a transformation $\underline{\textnormal{A}}=\underline{\textnormal{T}}\,\underline{\textnormal{$\Lambda$}}\,\underline{\textnormal{T}}^{T}$ so that
@@
where $\mathbf{W}=\underline{\textnormal{T}}^{T}\mathbf{U}$ defines a vector of (nonlinear) characteristic variables. The variables $\mathbf{U}$ come from the skew-symmetric analysis of the governing equation and may represent either primitive variables, as in the incompressible Euler equations, or Roe-type variables involving square roots, as in the shallow water equations [13]. The sign of the terms in the diagonal matrix $\underline{\textnormal{$\Lambda$}}$ specifies the minimal number of boundary conditions required to obtain a bound on the solution in terms of external data [33, 13, 34, 28]. For linear problems, $\underline{\textnormal{T}}$ is the matrix of eigenvectors and $\underline{\textnormal{$\Lambda$}}$ is the matrix of eigenvalues of the matrix $\underline{\textnormal{A}}$. For nonlinear problems we will use a congruence transformation of $\underline{\textnormal{A}}$ to determine an appropriate transformation $\underline{\textnormal{T}}$ that relates nonlinear and linear theory, as discussed in Section 6.1.
The decomposition in (4.1) exposes how solution information propagates in the directions determined by the sign of the terms in the diagonal matrix $\underline{\textnormal{$\Lambda$}}$, so we decompose $\underline{\textnormal{$\Lambda$}}=\underline{\textnormal{$\Lambda$}}^{+}+\underline{\textnormal{$\Lambda$}}^{-}$, where $\underline{\textnormal{$\Lambda$}}^{\pm}=\frac{1}{2}\left(\underline{\textnormal{$\Lambda$}}\pm|\underline{\textnormal{$\Lambda$}}|\right)$ into matrices with only positive or negative entries. Possible zero entries are included in $\underline{\textnormal{$\Lambda$}}^{+}$. We associate outgoing and incoming (nonlinear) characteristic variables with those positive and negative terms by
@@
Again, following [13, 28, 29], we prove the following purely algebraic result:
@@
$\mathbf{W}^{T}\underline{\textnormal{$\Lambda$}}\mathbf{W}+\mathbf{U}^{T}\left(2\underline{\textnormal{T}}\,\underline{\textnormal{I}}^{-}\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\left(\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\mathbf{W}^{-}-\mathbf{G}\right)\right)\geq-\mathbf{G}^{T}\mathbf{G},$ (4.3)
@@
Now let $\mathbf{G}$ consist of accurate boundary data in the form of the boundary operator $\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\mathbf{W}^{-}$ such that $\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\mathbf{W}^{-}-\mathbf{G}$ is small, and interpret $2\underline{\textnormal{T}}\,\underline{\textnormal{I}}^{-}\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\left(\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\mathbf{W}^{-}-\mathbf{G}\right)$ as the simultaneous approximation term (SAT) [18]. Then
@@
We compare the terms in (4.8) with (4.3) and (4.6) to arrive at the desired result (4.7). ∎
@@
Theorems 4.1 and 4.2 illustrate that $\mathbf{G}$ contains the external boundary data needed to balance the boundary operator $\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\mathbf{W}^{-}$ and acquire a bound. For complete details on the role of $\mathbf{G}$ in the nonlinear SATs see [13, 28, 29, 30]. For the Burgers equation and the shallow water equations specific forms of $\mathbf{G}$ and how they relate to the numerical flux functions will be derived in Secs. 5.1 and 6.2, respectively.
@@
Remark 4.4.
The relation (4.6) yields a valid numerical flux provided there exists a common set of variables, $\mathbf{Z}$, between the entropy variables $\mathbf{V}$ and the variables $\mathbf{U}$. The three variable sets constitute different representations of the same physical state and are related by invertible transformations. It is always possible to find $\mathbf{V}(\mathbf{Z})=\underline{\textnormal{M}}_{\mathbf{V}}(\mathbf{Z})\mathbf{Z}$ and $\mathbf{U}(\mathbf{Z})=\underline{\textnormal{M}}_{\mathbf{U}}(\mathbf{Z})\mathbf{Z}$ to rewrite the variables in terms of the common variables $\mathbf{Z}$ via an outer product representation [35], e.g.,
$\underline{\textnormal{M}}_{\mathbf{V}}(\mathbf{Z})=\frac{\mathbf{V}\mathbf{Z}^{T}}{\mathbf{Z}^{T}\mathbf{Z}}\quad\text{so that}\quad\underline{\textnormal{M}}_{\mathbf{V}}(\mathbf{Z})\mathbf{Z}=\frac{\mathbf{V}\mathbf{Z}^{T}}{\mathbf{Z}^{T}\mathbf{Z}}\mathbf{Z}=\mathbf{V}.$ (4.9)
Importantly, the matrices $\underline{\textnormal{M}}_{\mathbf{V}}(\mathbf{Z})$ and $\underline{\textnormal{M}}_{\mathbf{U}}(\mathbf{Z})$ are not unique and their entries may depend nonlinearly on $\mathbf{Z}$. Note that these matrices are not Jacobians; they describe the variable transformation itself, rather than its differential. Substituting the transformations, (4.6) becomes
$(\underline{\textnormal{M}}_{\mathbf{V}}\mathbf{Z})^{T}\left(\mathbf{F}^{*}_{n}-\mathbf{F}_{n}\right)=(\underline{\textnormal{M}}_{\mathbf{U}}\mathbf{Z})^{T}\left(2\underline{\textnormal{T}}\,\underline{\textnormal{I}}^{-}\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\left(\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\mathbf{W}^{-}-\mathbf{G}\right)\right).$ (4.10)
Manipulating this expression we find the provably stable numerical boundary flux function takes the form
$\mathbf{F}^{*}_{n}=2\underline{\textnormal{M}}_{\mathbf{V}}^{-T}\underline{\textnormal{M}}_{\mathbf{U}}^{T}\,\underline{\textnormal{T}}\,\underline{\textnormal{I}}^{-}\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\left(\sqrt{|\underline{\textnormal{$\Lambda$}}^{-}|}\mathbf{W}^{-}-\mathbf{G}\right)+\mathbf{F}_{n},$ (4.11)
provided that the characteristic-type variables $\mathbf{W}$ and diagonal matrix $\underline{\textnormal{$\Lambda$}}$ can be found.
In Sec. 6, we use the primitive variables as a common variable set $\mathbf{Z}$ for the two-dimensional shallow water equations. We then choose a diagonal matrix $\underline{\textnormal{$\Lambda$}}$ and describe how to create a transformation of the boundary matrix to reveal a set of characteristic-type variables. Then, nonlinear stable boundary fluxes can be determined. To summarize, once a choice for a common variable set is made, with corresponding transformations $\underline{\textnormal{M}}_{\mathbf{V}}(\mathbf{Z})$ and $\underline{\textnormal{M}}_{\mathbf{U}}(\mathbf{Z})$, and a suitable transformation matrix $\underline{\textnormal{T}}$ in the decomposition of the boundary matrix is created for a chosen diagonal matrix $\underline{\textnormal{$\Lambda$}}$, then the numerical flux is uniquely determined from (4.11).
@@ 5 A stable nonlinear inflow boundary flux for the Burgers equation
To illustrate the process and to motivate and highlight the vital steps, we demonstrate the use of Thm. 4.2 by deriving an energy stable boundary flux for the Burgers equation. Although nonlinear, the Burgers equation is sufficiently simple that the boundary flux function in (3.1) can be derived via the energy method and a nonstandard rewriting of the boundary terms. We show here that the derivation using the energy method gives the same result as Thm. 4.2.
5.1 Boundary flux derivation for Burgers’ equation
@@
When $u>0$, $\left.\frac{1}{3}u^{3}\right|_{x_{R}}>0$ and has the correct sign to dissipate energy, so no boundary condition is needed at $x_{R}$. At the left boundary, if we ensure that the contribution is a square of specified data, i.e., if the contribution from the left (inflow) side is equal to some $G^{2}$, where $G(t)$ is a function solely of specified data, then the energy is bounded exclusively by data, i.e., $\mathcal{E}_{t}\leq G^{2}$ and $\mathcal{E}(T)\leq\int_{0}^{T}G^{2}dt$.
@@
$G=\overline{W}(u_{\textrm{ext}})=\sqrt{\frac{1}{3}|u_{\textrm{ext}}|}u_{\textrm{ext}},$ (5.4)
then $BT\leq G^{2}$ where $G$ is a function only of boundary data. This means that, unlike in the linear analysis, one does not specify the solution at the boundary, but (5.4) instead, which includes the wave speed.
Since (3.1) imposes the boundary conditions in weak form as a penalty, we need the appropriate penalty that ensures that $\mathcal{E}$ is bounded by data in the same way. From (5.4), we penalize the PDE by adding a term proportional to $\overline{W}-G$, where $G$ is defined in (5.4),
@@
and $l(x)$ is a lifting operator defined as $\int_{x_{L}}^{x_{R}}\phi l(\psi)dx=\phi\psi|_{x_{L}}$ [36, 37]. The penalty is on the characteristic variable $\overline{W}$, which, again, differs from linear problems where the dependent variable $u$ is penalized towards the external solution, $u_{\textrm{ext}}$.
@@
When we substitute for $\overline{W}$, $G$, $F$, and $\sigma$, use the fact that $u>0$, and rearrange, we find the numerical boundary flux function that satisfies Thm. 4.2 and ensures that the entropy is bounded only by data to be
@@
This boundary term is not a square bounded by data independent of $u$, and hence does not lead to an estimate solely in terms of data. If, instead, we use the local Lax-Friedrichs (LLF) flux [38] at the boundary, then
$BT=\left\{\begin{aligned} &-\frac{u^{3}}{6}+\frac{uu_{\textrm{ext}}^{2}}{2}-\frac{u(u-u_{\textrm{ext}})^{2}}{4},\quad u>u_{\textrm{ext}}\\[5.69046pt] &-\frac{u^{3}}{6}+\frac{uu_{\textrm{ext}}^{2}}{2}+\frac{u(u-u_{\textrm{ext}})^{2}}{4},\quad u<u_{\textrm{ext}}.\end{aligned}\right$. (5.13)
Again, the boundary term is not bounded independently of the interior state, $u$, and using the flux (5.13) does not lead to an entropy bound.